# Sharp Time Decay and Regularization

### v1

@@ 3.1 Existence and uniqueness
[37, Theorem 5.10] Let $m>0$ and assume (G1). Then, for every initial datum $u_{0}\in L^{1+m}(\Omega)\cap H^{-1}(\Omega)$ there exits a unique $H^{*}$ solution $u=u(t,x)$ of (CPME). Moreover, it satisfies that
@@
Convergence of the discrete scheme in a more general setting has been proven very recently by Li and Salgado in. [38, Theorem 4.5.].
@@
$\begin{split}u\in L^{\infty}((0,T):H^{*}(\Omega)),\quad|u|^{m-1}u\in L^{2}((0,T):H(\Omega))\,,\\ \ell*\left\|D^{\alpha}_{t}u\right\|_{H^{*}(\Omega)}^{2}\in L^{\infty}(0,T)\qquad\qquad\,,\end{split}$ (29)
@@
Since there is uniqueness of $H^{*}$-solution, Theorem 3.1 and Theorem 3.2 describe the same solution for a fixed $u_{0}\in L^{1+m}(\Omega)\cap H^{*}(\Omega)$. Despite the difference in the approximation procedures, the limit solution coincide. Hence, we can use the estimates from both theorems simultaneously.
@@ 3.2 Comparison Principle
Now, we are in a position to prove the comparison principle for the $H^{*}$-solutions of (CPME). Let us recall the statement which holds for every $m>0$.
@@ 4 Smoothing effects
The aim of this section is to prove the $L^{p}-L^{\infty}$ smoothing effects in Theorem 1.4. We divide the proof in several steps. First, in Proposition 4.2 we use the dual formulation of the problem and the Fundamental Theorem of Calculus for Caputo derivative to obtain a fundamental pointwise formula. Then, in Theorem 4.3 we obtain a time monotonicity formulae for the function $t^{\frac{\alpha}{m-1}}u(t)$. The time monotonicity formulae we obtain are consistent with the case $\alpha=1$, since they depend only on the time scaling and the comparison principle. Next, we apply the time monotonicity formulae in the fundamental pointwise formula in order to obtain the upper bounds of $u^{m}(t,x)$, as described in Proposition 4.4. We conclude the proof of Theorem 1.4 and Theorem 4.6 using assumptions (G1) and (G2), respectively, and Hölder’s inequality.
@@
We obtain that $f*g\in L^{2}((0,T))$ by Youngs inequality. For any $\varphi\in C_{c}^{\infty}([0,T])$ let us consider
@@
Taking in to account estimate (29) from the Existence Theorem 3.2, we make the dual formulation of the problem rigourous and we prove the Fundamental Theorem of Calculus for Caputo in the following proposition.
@@
*$\ell*\|D^{\alpha}_{t}u\|^{2}_{H^{*}(\Omega)}\in L^{\infty}(\Omega)\,.$ (39)
@@
$\int_{0}^{T}\int_{\Omega}D^{\alpha}_{t}u(\tau,x)\mathcal{L}^{-1}\psi(\tau,x)\,{\rm d}x\,{\rm d}\tau$ =*$=\ell*\left(\int_{\Omega}D^{\alpha}_{t}u\mathcal{L}^{-1}\delta_{x_{0}}\,{\rm d}x\right)=\ell*k*\left(\frac{\rm d}{\,{\rm d}t}\int_{\Omega}(u(t,x)\mathbb{G}_{\Omega}(x_{0},x)\,{\rm d}x\right)$
@@
Note that condition (39) implies that $\tilde{\psi}$ is an admissible test function, and *$\ell*\left(D^{\alpha}_{t}u\mathcal{L}^{-1}\varphi\,{\rm d}x\right)\in L^{\infty}(0,T)$, since
@@
*$\ell*\left(\int_{\Omega}D^{\alpha}_{t}u\mathcal{L}^{-1}\varphi\,{\rm d}x\right)\!(t)\!$ $=\!\!\int_{\Omega}\!\!\ell*\!\left(\frac{\rm d}{\,{\rm d}t}[k*(u-u_{0})]\right)\!\!(t)\mathcal{L}^{-1}\varphi\,{\rm d}x=\!\int_{\Omega}\!\!\frac{\rm d}{\,{\rm d}t}\left(\ell*[k*(u-u_{0})](t)\right)\mathcal{L}^{-1}\varphi\,{\rm d}x$
@@
As in the previuos case, we use comparison principle of Theorem 1.2 in order to obtain $u_{\lambda}(t,x)\leq u(t,x)$ for any $t>0$ and a.e. $x\in\Omega$. Therefore, we choose $t=t_{1}$ to get
@@
The idea of the proof is to use Benilan-Crandall estimates (4.3) on the right hand side of the fundamental integral formula (40). Namely, it sufices to prove the following inequality
@@
and consider the nonnegative $H^{*}$-solutions $\tilde{u}^{+}$ and $\tilde{u}^{-}$ with nonnegative initial data $(u_{0})_{+},(u_{0})_{-}$, respectively. First, we have to check that $(u_{0})_{+},(u_{0})_{-}\in L^{p}(\Omega)\cap H^{*}(\Omega)$ with $p$ as in (5). Since $u_{0}\in L^{p}(\Omega)\cap H^{*}(\Omega)$ with $p$ satisfying (5), then $(u_{0})_{+},(u_{0})_{-}\in L^{p}(\Omega)$ with the same $p$. It remains to prove that they also belong to $H^{*}(\Omega)$. For this purpose, we have to use the Hardy-Littlewood-Sobolev inequality associated to the quadratic form of the operator: There exists $\mathcal{H}_{\mathcal{L}}>0$ such that
@@
which holds under assumption (G1), see [15, Theorem 7.5]. Since $N>2s$ then $\frac{2N}{N+2s}<\frac{N}{2s}$, and hence
@@ 5 Sharp time decay of $L^{p}$-norms
We first show, heuristically, why we should not expect extinction in finite time as in the case with classical time derivative. Let us consider solutions of for (CPME) of the form
@@
*$\ell*\|D^{\alpha}_{t}u(t)\|^{2}_{H^{*}(\Omega)}\in L^{\infty}((0,T))\,,$
@@
*$\ell*\|D^{\alpha}_{t}u(t)\|_{L^{1}_{\Phi_{1}}(\Omega)}\in L^{\infty}((0,T))\,.$ (52)
@@
|*$\left|\ell*\left(\int_{\Omega}D^{\alpha}_{t}u(t,x)\Phi_{1}(x)\,{\rm d}x\right)\right|$ $=\left|\int_{0}^{t}\int_{\Omega}\ell(t-s)D^{\alpha}_{t}u(s,x)\Phi_{1}(x)\,{\rm d}x\,{\rm d}t\right|$
@@
where in the second inequality we have used Young’s inequality. We conclude by taking the essential supremum in $t\in(0,T)$ and using estimate (29). ∎
The last lemma is used in the proof of a lower and upper bounds for the $L^{1}_{\Phi_{1}}$-norm of the $H^{*}$-solutions.
@@
*$\ell*\left(\int_{\Omega}\partial_{t}[k*(u-u_{0})](t)\Phi_{1}\,{\rm d}x\right)=\int_{\Omega}\left(\ell*\partial_{t}[k*(u-u_{0})]\right)(t)\Phi_{1}\,{\rm d}x\leq 0\,.$
Since *$\ell*k=1$ and $\partial_{t}\left[1*(u(t,x)-u_{0}(x))\right]=u(t,x)-u_{0}(x)$ for a.e. $x\in\Omega$, we conclude that
@@
Proof of Theorem 1.7. The proof follows by an interpolation argument, using the lower bound of the $L^{1}_{\Phi_{1}}$-norm of solutions and the $L^{p}-L^{\infty}$ smoothing effects. By Proposition 5.2 and Theorem 1.4 we obtain that for every $t>0$
@@ 6 Open questions
    Asymptotic behaviour: Theorem 1.7 shows that solutions of (CPME) does not vanish in finite time for any $m>0$. One may wonder about the behaviour of the solutions when $t\rightarrow\infty$. From the smoothing effects of Theorem 1.4, we deduce that solution vanish at infinity, but we would like to know if there is any asymptotic profile $S(x)$ such that
@@
    for certain norm $X$. From the classical theory, we expect that $S$ is the solution of the elliptic problem described in Section 5 and that the solution behaves at infinity like the separate-variable solution. In the classical approach, the results relies on a change of variables which with Caputo time derivative is more involved, see the fundamental identity for regular kernels in Lemma A.4.
@@
In this appendix we collect some of the known results about the Caputo derivative and the associated fractional ODEs that are useful for our study. Let us start with the linear differential equation with $\alpha\in(0,1)$, $v_{0}>0$ and $\lambda\in\mathbb{R}^{N}$
@@
[35, Example 4.11] Let $10<\alpha<$ 1 and $\beta>-\alpha$. Then, the solution of
@@
[19, Proposition 6] Let $10<\alpha<$ 1, $r>0$ and $l>r-\frac{1}{\alpha}$, then
[…]
@@
[49, Theorem 7.1] Let $\alpha\in(0,1)$, $\lambda,m>0$ and $v_{0}>0$. Let $v\in H^{1}_{\rm loc}(\mathbb{R}_{+})$ be the solution of
@@
The explicit form of the solution to this basic nonlinear fractional ODE is not known, but in [49] they construct subsolution and supersolution to estimate the lower and upper bounds. Note that the behaviour of $v$ described in (64) is completely different from the solution of the analogous classical ODE. In the classical case, the exponent $m$ plays a crucial role. Namely, for the classical ODE of the form $\partial_{t}v=-\lambda v^{m}$ with $v(0)=v_{0}$ we have that:
@@
This shows at the level of ODEs how different are the solutions when we consider Caputo time derivative. In particular, there is no continuity of results when $\alpha\rightarrow 1$.
@@
[49, Lemma 2.2] Let $T>0$ and $I$ an open set of $\mathbb{R}$. Further, let $\tilde{k}\in W^{1,1}([0,T])$, $h\in C^{1}(I)$ and $u\in L^{1}((0,T))$ with $u(t)\in I$ for a.e. $t\in(0,T)$. Suppose that the functions $h(u),h^{\prime}(u)u$ and $h^{\prime}(u)(\tilde{k}^{\prime}*u)$ belong to $L^{1}((0,T))$ (which is the case if $u\in L^{\infty}((0,T))$). Then, we have for a.e. $t\in(0,T)$

### v3

@@ 3.1 Existence and uniqueness
[54, Theorem 5.10] Let $m>0$ and assume (G1). Then, for every initial datum $u_{0}\in L^{1+m}(\Omega)\cap H^{-1}(\Omega)$ there exits a unique $H^{*}$ solution $u=u(t,x)$ of (CPME). Moreover, it satisfies that
@@
Convergence of the discrete scheme in a more general setting has been proven very recently by Li and Salgado in. [55, Theorem 4.5.].
@@
$\begin{split}u\in L^{\infty}((0,T):H^{*}(\Omega)),\quad|u|^{m-1}u\in L^{2}((0,T):H(\Omega))\,,\\ \ell*\left\|D^{\alpha}_{t}u\right\|_{H^{*}(\Omega)}^{2}\in L^{\infty}((0,T))\qquad\qquad\,,\end{split}$ (38)
@@
Since there is uniqueness of $H^{*}$-solution, Theorem 3.1 and Theorem 3.2 describe the same solution for a fixed $u_{0}\in L^{1+m}(\Omega)\cap H^{*}(\Omega)$. Despite the difference in the approximation procedures, the limit solution coincides. Hence, we can use the estimates from both theorems simultaneously.
@@ 3.2 Comparison Principle
Now, we are ready to prove the comparison principle for the $H^{*}$-solutions of (CPME). Let us recall the precise statement:
@@ 4 Smoothing effects
The aim of this section is to prove the $L^{p}-L^{\infty}$ smoothing effects stated in Theorem 1.4. We divide the proof in several steps. First, in Proposition 4.2 we use the dual formulation of the problem and the Fundamental Theorem of Calculus for Caputo derivative to obtain a fundamental pointwise formula. Then, in Theorem 4.3 we obtain a time monotonicity formulae for the function $t^{\frac{\alpha}{m-1}}u(t)$. The time monotonicity formulae we obtain are consistent with the case $\alpha=1$, since they depend only on the time scaling and the comparison principle. Next, we apply the time monotonicity formulae in the fundamental pointwise formula in order to obtain the upper bounds of $u^{m}(t,x)$, as described in Proposition 4.4. We conclude the proof of Theorem 1.4 and Theorem 4.6 using assumptions (G1) and (G2), respectively, and Hölder’s inequality.
@@
We obtain that $f*g\in L^{2}((0,T))$ by Young’s inequality. For any $\varphi\in C_{c}^{\infty}([0,T])$ let us consider
@@
Taking in to account estimate (38) from the Existence Theorem 3.2, we make the dual formulation of the problem rigorous and we prove the Fundamental Theorem of Calculus for Caputo in the following proposition.
@@
$\ell*\|D^{\alpha}_{t}u\|^{2}_{H^{*}(\Omega)}\in L^{\infty}((0,T))\,.$ (48)
@@
$\int_{0}^{T}\int_{\Omega}D^{\alpha}_{t}u(\tau,x)\mathcal{L}^{-1}\psi(\tau,x)\,{\rm d}x\,{\rm d}\tau$ $=\ell*\left(\int_{\Omega}D^{\alpha}_{t}u\mathcal{L}^{-1}\delta_{x_{0}}\,{\rm d}x\right)=\ell*k*\left(\frac{\rm d}{\,{\rm d}t}\int_{\Omega}(u(t,x)\mathbb{G}_{\Omega}(x_{0},x)\,{\rm d}x\right)$
@@
Note that condition (48) implies that $\tilde{\psi}$ is an admissible test function, and $\ell*\left(D^{\alpha}_{t}u\mathcal{L}^{-1}\varphi\,{\rm d}x\right)\in L^{\infty}(0,T)$, since
@@
$\ell*\left(\int_{\Omega}D^{\alpha}_{t}u\mathcal{L}^{-1}\varphi\,{\rm d}x\right)\!(t)\!$ $=\!\!\int_{\Omega}\!\!\ell*\!\left(\frac{\rm d}{\,{\rm d}t}[k*(u-u_{0})]\right)\!\!(t)\mathcal{L}^{-1}\varphi\,{\rm d}x=\!\int_{\Omega}\!\!\frac{\rm d}{\,{\rm d}t}\left(\ell*[k*(u-u_{0})](t)\right)\mathcal{L}^{-1}\varphi\,{\rm d}x$
@@
As in the previous case, we use comparison principle of Theorem 1.2 in order to obtain $u_{\lambda}(t,x)\leq u(t,x)$ for any $t>0$ and a.e. $x\in\Omega$. Therefore, we choose $t=t_{1}$ to get
@@
The idea of the proof is to use Benilan-Crandall estimates (4.3) on the right hand side of the fundamental integral formula (49). Namely, it suffices to prove the following inequality
@@
and consider the nonnegative $H^{*}$-solutions $\tilde{u}^{+}$ and $\tilde{u}^{-}$ with nonnegative initial data $(u_{0})_{+},(u_{0})_{-}$, respectively. First, we have to check that $(u_{0})_{+},(u_{0})_{-}\in L^{p}(\Omega)\cap H^{*}(\Omega)$ with $p$ as in (5). Since $u_{0}\in L^{p}(\Omega)\cap H^{*}(\Omega)$ with $p$ satisfying (5), then $(u_{0})_{+},(u_{0})_{-}\in L^{p}(\Omega)$ with the same $p$. It remains to prove that they also belong to $H^{*}(\Omega)$. For this purpose, we use the Hardy-Littlewood-Sobolev inequality associated to the quadratic form of the operator: There exists $\mathcal{H}_{\mathcal{L}}>0$ such that
@@
which holds under assumption (G1), see [18, Theorem 7.5]. Since $N>2s$ then $\frac{2N}{N+2s}<\frac{N}{2s}$, and hence
@@ 5 Sharp time decay of $L^{p}$-norms
Let us explain, heuristically, why one should not expect extinction in finite time, unlike the case with classical time derivative. Let us consider solutions of for (CPME) of the form
@@
$\ell*\|D^{\alpha}_{t}u(t)\|^{2}_{H^{*}(\Omega)}\in L^{\infty}((0,T))\,,$
@@
$\ell*\|D^{\alpha}_{t}u(t)\|_{L^{1}_{\Phi_{1}}(\Omega)}\in L^{\infty}((0,T))\,.$ (61)
@@
$\left|\ell*\left(\int_{\Omega}D^{\alpha}_{t}u(t,x)\Phi_{1}(x)\,{\rm d}x\right)\right|$ $=\left|\int_{0}^{t}\int_{\Omega}\ell(t-s)D^{\alpha}_{t}u(s,x)\Phi_{1}(x)\,{\rm d}x\,{\rm d}t\right|$
@@
thanks to Young’s inequality in the third line. We conclude by taking the essential supremum in $t\in(0,T)$ and using estimate (38). ∎
The last lemma provides a lower and upper bounds for the $L^{1}_{\Phi_{1}}$-norm of the $H^{*}$-solutions.
@@
$\ell*\left(\int_{\Omega}\partial_{t}[k*(u-u_{0})](t)\Phi_{1}\,{\rm d}x\right)=\int_{\Omega}\left(\ell*\partial_{t}[k*(u-u_{0})]\right)(t)\Phi_{1}\,{\rm d}x\leq 0\,.$
Since $\ell*k=1$ and $\partial_{t}\left[1*(u(t,x)-u_{0}(x))\right]=u(t,x)-u_{0}(x)$ for a.e. $x\in\Omega$, we conclude that
@@
Proof of Theorem 1.7. The proof follows by an interpolation argument, the lower bound of the $L^{1}_{\Phi_{1}}$-norm, and the $L^{p}-L^{\infty}$ smoothing effects. By Proposition 5.2 and Theorem 1.4 we obtain that for every $t>0$
@@ 6 Open questions
    Asymptotic behaviour: Theorem 1.7 shows that solutions of (CPME) does not vanish in finite time for any $m>0$. One may wonder about the behaviour of the solutions when $t\rightarrow\infty$. From the smoothing effects of Theorem 1.4, we deduce that solutions vanish at time increases, but we would like to know if there is any asymptotic profile $S(x)$ such that
@@
    for certain norm $X$. From the classical theory, we expect that $S$ is the solution of the elliptic problem described in Section 5 and that the solution behaves at infinity like the the separate-variable solution. In the classical approach, the results relies on a change of variables; with Caputo time derivative this approach would be more complicated, see the fundamental identity for regular kernels in Lemma A.4.
@@
We collect some known results about the Caputo derivative and the associated fractional ODEs that are useful for our study. Let us start with the linear differential equation with $\alpha\in(0,1)$, $v_{0}>0$ and $\lambda\in\mathbb{R}^{N}$
@@
[51, Example 4.11] Let $10<\alpha<$ 1 and $\beta>-\alpha$. Then, the solution of
@@
[22, Proposition 6] Let $10<\alpha<$ 1, $r>0$ and $l>r-\frac{1}{\alpha}$, then
$E_{\alpha,r,l}(-x)\asymp\frac{1}{1+c\,x}\qquad\quad\mbox{as}\quad x\rightarrow\infty\,.$
@@
[73, Theorem 7.1] Let $\alpha\in(0,1)$, $\lambda,m>0$ and $v_{0}>0$. Let $v\in H^{1}_{\rm loc}(\mathbb{R}_{+})$ be the solution of
@@
The explicit form of the solution to this basic nonlinear fractional ODE is not known, but in [73] the authors construct subsolution and supersolution to estimate the lower and upper bounds. Note that the behaviour of $v$ described in (73) is completely different from the solution of the analogous classical ODE. In the classical case, the value of the exponent $m$ plays a crucial role. Namely, for the classical ODE of the form $\partial_{t}v=-\lambda v^{m}$ with $v(0)=v_{0}$ we have that:
@@
Notice, at the level of ODEs, the difference of the asymptotic profile between classical and fractional time derivative. Moreover, there is no continuity of results when $\alpha\rightarrow 1$.
@@
[73, Lemma 2.2] Let $T>0$ and $I$ an open set of $\mathbb{R}$. Further, let $\tilde{k}\in W^{1,1}([0,T])$, $h\in C^{1}(I)$ and $u\in L^{1}((0,T))$ with $u(t)\in I$ for a.e. $t\in(0,T)$. Suppose that the functions $h(u),h^{\prime}(u)u$ and $h^{\prime}(u)(\tilde{k}^{\prime}*u)$ belong to $L^{1}((0,T))$ (which is the case if $u\in L^{\infty}((0,T))$). Then, we have for a.e. $t\in(0,T)$